OpenStudy (anonymous):

A collection of quarters and nickels is worth $14.80. There are 88 coins in all. How many of each coin are there? ____ quarters ____ nickels @jojo4eva

OpenStudy (anonymous):

x = (1480-880)/15 = 40, the number of quarters. 48 Dimes And...checking 1000 + 480 = 1480 0r $14.80 So 40 quarters And 48 Dimes

OpenStudy (anonymous):

Thank you..... MEDAL TIME

OpenStudy (anonymous):

There ya go